# Keldysh field theory of dynamical exciton condensation transitions in nonequilibrium electron-hole bilayers

### v2

@@ I.1 Keldysh formalism
$S_{\rm inter}[\bar{\Delta},\Delta,\bar{\psi},\psi]=-\frac{1}{A}\sum_{\tau\tau^{\prime}}\int_{C}dt_{C}\left\{\sum_{kk^{\prime}q}V_{kk^{\prime}}^{-1}\bar{\Delta}_{kq}^{\tau\tau^{\prime}}(t_{C})\Delta_{k^{\prime}q}^{\tau\tau^{\prime}}(t_{C})+\sum_{kq}\big{[}\Delta_{kq}(t_{C})\bar{\psi}_{\tau c,k+q/2}(t_{C})\psi_{\tau^{\prime}v,k^{\prime}-q/2}(t_{C})+{\rm c.c.}\big{]}\right\},$ (11)
where $V_{kk^{\prime}}^{-1}$ is the $(k,k^{\prime})$-component of the inverse of the interlayer Coulomb interaction matrix $V(k-k^{\prime})$. After the Keldysh rotation (‘c’ and ‘q’ stand for ‘classical’ and ‘quantum’ respectively)
@@
$\begin{split}S_{\rm inter}&=-\frac{1}{A}\sum_{\tau\tau^{\prime}}\sum_{kk^{\prime}q}\int\frac{d\omega}{2\pi}\,V_{kk^{\prime}}^{-1}\bar{\Delta}_{kq}^{\tau\tau^{\prime}}(\omega)\alpha_{1}\Delta_{k^{\prime}q}^{\tau\tau^{\prime}}(\omega)\\ &-\frac{1}{\sqrt{2}A}\sum_{\tau\tau^{\prime}}\sum_{kq}\int\frac{d\omega}{2\pi}\int\frac{d\nu}{2\pi}\left\{\bar{\psi}_{\tau c,k+q/2}(\nu+\frac{\omega}{2})\big{[}\Delta_{kq}^{{\rm c},\tau\tau^{\prime}}(\omega)\alpha_{0}+\Delta_{kq}^{{\rm q},\tau\tau^{\prime}}(\omega)\alpha_{1}\big{]}\psi_{\tau v,k-q/2}(\nu-\frac{\omega}{2})+{\rm c.c.}\right\},\end{split}$ (13)
@@
Figure S1: Diagrammatic representations of the coefficients of (a) $\bar{\Delta}^{\rm c}\Delta^{\rm q}$; (b) $\bar{\Delta}^{\rm q}\Delta^{\rm q}$; (c) $\Delta^{\rm c}\Delta^{\rm q}$; (d) $\Delta^{\rm q}\Delta^{\rm q}$. Here the dotted lines represent the bosonic fields $\bar{\Delta},\Delta$ (or the $\bar{\Phi},\Phi$ fields after projecting onto the 1s-exciton basis), solid lines represent Green functions for electrons (Eqs. I.1), and the crosses represent interlayer tunneling.
@@ I.2 Classical-quantum quadratic Green function
$\frac{i}{2A^{2}}\sum_{\tau\tau^{\prime}}\sum_{kq}\int\frac{d\omega}{2\pi}\bar{\Delta}_{kq}^{{\rm c},\tau\tau^{\prime}}(\omega)\Delta_{kq}^{{\rm q},\tau\tau^{\prime}}(\omega)\int\frac{d\nu}{2\pi}\left[G_{c,k+q/2}^{K}(\nu+\frac{\omega}{2})G_{v,k-q/2}^{R}(\nu-\frac{\omega}{2})+G_{c,k+q/2}^{A}(\nu+\frac{\omega}{2})G_{v,k-q/2}^{K}(\nu-\frac{\omega}{2})\right].$ (14)
@@
$S_{2,0}^{\rm cq}[\bar{\Delta},\Delta]=-\frac{1}{A^{2}}\sum_{\tau\tau^{\prime}}\sum_{kq}\int\frac{d\omega}{2\pi}\,\bar{\Delta}_{kq}^{{\rm q},\tau\tau^{\prime}}(\omega)\left[\frac{1}{\omega-\xi_{k+q/2}-\xi_{k-q/2}}\Delta_{kq}^{{\rm c},\tau\tau^{\prime}}(\omega)+A\sum_{k^{\prime}}V_{kk^{\prime}}^{-1}\Delta_{k^{\prime}q}^{{\rm c},\tau\tau^{\prime}}(\omega)\right]+{\rm c.c.}$ (15)
@@
$\frac{k^{2}}{m}\varphi_{k}-\frac{1}{A}\sum_{k^{\prime}}V_{kk^{\prime}}\varphi_{k^{\prime}}=-E_{b}\varphi_{k},$ (16)
where $E_{b}$ is the binding energy of the 1s exciton mode, $S_{2,0}^{\rm cq}$ is equivalently expressed in terms of the $\bar{\Phi},\Phi$ fields as
$S_{2,0}^{\rm cq}[\bar{\Phi},\Phi]=\frac{1}{A}\sum_{\tau\tau^{\prime}q}\int\frac{d\omega}{2\pi}(\omega-\frac{q^{2}}{4m}-E_{g}+E_{b}+i\gamma)\bar{\Phi}_{q}^{{\rm q},\tau\tau^{\prime}}(\omega)\Phi_{q}^{{\rm c},\tau\tau^{\prime}}(\omega)+{\rm c.c.},$ (17)
in the limit $\omega\approx E_{g}-E_{b}+q^{2}/4m$. An imaginary coefficient $i\gamma$ that describes coupling to the reservoirs is included for completeness. Lowest-order expansion in $\Gamma_{c,v}$ gives $\gamma\propto\Gamma_{c}\Gamma_{v}(\omega-\mu_{c}+\mu_{v})$.
@@ I.3 Quantum-quantum quadratic Keldysh term
$\frac{i}{2}\sum_{\rm qq}\int GG=\frac{2i\Gamma_{c}\Gamma_{v}/\pi}{(\xi_{+}-\mu_{c})^{2}(\xi_{-}+\mu_{v})^{2}}(\omega-\mu_{x})\coth\frac{\omega-\mu_{x}}{2T},$ (19)
where $\xi_{\pm}=\xi_{k\pm q/2}$ and $\mu_{x}=\mu_{c}-\mu_{v}$ is the exciton chemical potential. The quantum-quantum quadratic action then reads
@@ I.4 Classical-quantum Josephson term
$\frac{i}{2}\sum_{\rm cq}\int GtGGtG=-\frac{\xi_{+}+\xi_{-}}{2\xi_{+}\xi_{-}[\omega^{2}-(\xi_{+}+\xi_{-})^{2}]}t_{\tau+}^{*}t_{\tau^{\prime}-}^{*}.$ (22)
@@
As an order-of-magnitude estimate, $c_{J}\sim mv_{t}^{2}E_{b}^{2}/E_{g}^{2}$.
@@ II.1 Large bias limit
$S_{0}[\bar{\Phi},\Phi]=\frac{1}{A}\sum_{q}\int\frac{d\omega}{2\pi}\,{\rm Tr}\bigg{[}(\omega-\frac{q^{2}}{4m}-E_{g}+E_{b}+i\gamma)\Phi^{\rm q\dagger}_{q}(\omega)\Phi^{\rm c}_{q}(\omega)+{\rm c.c.}+ig(\omega-\mu_{x})\coth\frac{\omega-\mu_{x}}{2T}\Phi^{\rm q\dagger}_{q}(\omega)\Phi^{\rm q}_{q}(\omega)\bigg{]}.$ (27)

### v3

@@ I.1 Keldysh formalism
$S_{\rm inter}[\bar{\Delta},\Delta,\bar{\psi},\psi]=-\frac{1}{A}\sum_{\tau\tau^{\prime}}\int_{C}dt_{C}\left\{\sum_{kk^{\prime}q}V_{kk^{\prime}}^{-1}\bar{\Delta}_{kq}^{\tau\tau^{\prime}}(t_{C})\Delta_{k^{\prime}q}^{\tau\tau^{\prime}}(t_{C})+\sum_{kq}\big{[}\Delta_{kq}(t_{C})\bar{\psi}_{\tau c,k+\frac{m_{e}^{*}}{M}q}(t_{C})\psi_{\tau^{\prime}v,k^{\prime}-\frac{m_{h}^{*}}{M}q}(t_{C})+{\rm c.c.}\big{]}\right\},$ (11)
where $M=m_{e}^{*}+m_{h}^{*}$ and $V_{kk^{\prime}}^{-1}$ is the $(k,k^{\prime})$-component of the inverse of the interlayer Coulomb interaction matrix $V(k-k^{\prime})$. After the Keldysh rotation (‘c’ and ‘q’ stand for ‘classical’ and ‘quantum’ respectively)
@@
$\begin{split}S_{\rm inter}&=-\frac{1}{A}\sum_{\tau\tau^{\prime}}\sum_{kk^{\prime}q}\int\frac{d\omega}{2\pi}\,V_{kk^{\prime}}^{-1}\bar{\Delta}_{kq}^{\tau\tau^{\prime}}(\omega)\alpha_{1}\Delta_{k^{\prime}q}^{\tau\tau^{\prime}}(\omega)\\ &-\frac{1}{\sqrt{2}A}\sum_{\tau\tau^{\prime}}\sum_{kq}\int\frac{d\omega}{2\pi}\int\frac{d\nu}{2\pi}\left\{\bar{\psi}_{\tau c,k+\frac{m_{e}^{*}}{M}q}(\nu+\frac{\omega}{2})\big{[}\Delta_{kq}^{{\rm c},\tau\tau^{\prime}}(\omega)\alpha_{0}+\Delta_{kq}^{{\rm q},\tau\tau^{\prime}}(\omega)\alpha_{1}\big{]}\psi_{\tau v,k-\frac{m_{h}^{*}}{M}q}(\nu-\frac{\omega}{2})+{\rm c.c.}\right\},\end{split}$ (13)
@@
Figure S1: Diagrammatic representations of the coefficients of (a) $\bar{\Delta}^{\rm c}\Delta^{\rm q}$; (b) $\bar{\Delta}^{\rm q}\Delta^{\rm q}$; (c) $\Delta^{\rm c}\Delta^{\rm q}$; (d) $\Delta^{\rm q}\Delta^{\rm q}$. Here the dotted lines represent the bosonic fields $\bar{\Delta},\Delta$ (or the $\bar{\Phi},\Phi$ fields after projecting onto the 1s-exciton basis), solid lines represent Green functions for electrons (Eqs. 7), and the crosses represent interlayer tunneling.
@@ I.2 Classical-quantum quadratic Green function
$\frac{i}{2A^{2}}\sum_{\tau\tau^{\prime}}\sum_{kq}\int\frac{d\omega}{2\pi}\bar{\Delta}_{kq}^{{\rm c},\tau\tau^{\prime}}(\omega)\Delta_{kq}^{{\rm q},\tau\tau^{\prime}}(\omega)\int\frac{d\nu}{2\pi}\left[G_{c,k+\frac{m_{e}^{*}}{M}q}^{K}(\nu+\frac{\omega}{2})G_{v,k-\frac{m_{h}^{*}}{M}q}^{R}(\nu-\frac{\omega}{2})+G_{c,k+\frac{m_{e}^{*}}{M}q}^{A}(\nu+\frac{\omega}{2})G_{v,k-\frac{m_{h}^{*}}{M}q}^{K}(\nu-\frac{\omega}{2})\right].$ (14)
@@
$S_{2,0}^{\rm cq}[\bar{\Delta},\Delta]=-\frac{1}{A^{2}}\sum_{\tau\tau^{\prime}}\sum_{kq}\int\frac{d\omega}{2\pi}\,\bar{\Delta}_{kq}^{{\rm q},\tau\tau^{\prime}}(\omega)\left[\frac{1}{\omega-(\xi_{c,k+\frac{m_{e}^{*}}{M}q}-\xi_{v,k-\frac{m_{h}^{*}}{M}q})}\Delta_{kq}^{{\rm c},\tau\tau^{\prime}}(\omega)+A\sum_{k^{\prime}}V_{kk^{\prime}}^{-1}\Delta_{k^{\prime}q}^{{\rm c},\tau\tau^{\prime}}(\omega)\right]+{\rm c.c.}$ (15)
@@
$\frac{k^{2}}{2m}\varphi_{k}-\frac{1}{A}\sum_{k^{\prime}}V_{kk^{\prime}}\varphi_{k^{\prime}}=-E_{b}\varphi_{k},$ (16)
where $m=m_{e}^{*}m_{h}^{*}/M$ is the reduced mass and $E_{b}$ is the binding energy of the 1s exciton mode, $S_{2,0}^{\rm cq}$ is equivalently expressed in terms of the $\bar{\Phi},\Phi$ fields as
$S_{2,0}^{\rm cq}[\bar{\Phi},\Phi]=\frac{1}{A}\sum_{\tau\tau^{\prime}q}\int\frac{d\omega}{2\pi}(\omega-\frac{q^{2}}{2M}-E_{g}+E_{b}+i\gamma)\bar{\Phi}_{q}^{{\rm q},\tau\tau^{\prime}}(\omega)\Phi_{q}^{{\rm c},\tau\tau^{\prime}}(\omega)+{\rm c.c.},$ (17)
in the limit $\omega\approx E_{g}-E_{b}+q^{2}/2M$. An imaginary coefficient $i\gamma$ that describes coupling to the reservoirs is included for completeness. Lowest-order expansion in $\Gamma_{c,v}$ gives $\gamma\propto\Gamma_{c}\Gamma_{v}(\omega-\mu_{c}+\mu_{v})$.
@@ I.3 Quantum-quantum quadratic Keldysh term
$\frac{i}{2}\sum_{\rm qq}\int GG=\frac{2i\Gamma_{c}\Gamma_{v}/\pi}{(\xi_{c+}-\mu_{c})^{2}(\xi_{v-}-\mu_{v})^{2}}(\omega-\mu_{x})\coth\frac{\omega-\mu_{x}}{2T},$ (19)
where $\xi_{b,\pm}=\xi_{b,k\pm q/2}$ and $\mu_{x}=\mu_{c}-\mu_{v}$ is the exciton chemical potential. The quantum-quantum quadratic action then reads
@@ I.4 Classical-quantum Josephson term
$\frac{i}{2}\sum_{\rm cq}\int GtGGtG=-\frac{1}{(\xi_{c+}-\xi_{v+})(\xi_{c-}-\xi_{v-})}\left(\frac{1}{\omega-\xi_{c+}+\xi_{v-}}-\frac{1}{\omega+\xi_{c-}-\xi_{v+}}\right)t_{\tau+}^{*}t_{\tau^{\prime}-}^{*}.$ (22)
@@
As an order-of-magnitude estimate, $c_{J}\sim mv_{t}^{2}E_{b}^{2}/E_{g}^{2}$ when $\omega\approx E_{g}-E_{b}$.
@@ II.1 Large bias limit
$S_{0}[\bar{\Phi},\Phi]=\frac{1}{A}\sum_{q}\int\frac{d\omega}{2\pi}\,{\rm Tr}\bigg{[}(\omega-\frac{q^{2}}{2M}-E_{g}+E_{b}+i\gamma)\Phi^{\rm q\dagger}_{q}(\omega)\Phi^{\rm c}_{q}(\omega)+{\rm c.c.}+ig(\omega-\mu_{x})\coth\frac{\omega-\mu_{x}}{2T}\Phi^{\rm q\dagger}_{q}(\omega)\Phi^{\rm q}_{q}(\omega)\bigg{]}.$ (27)